solve 3x-5)+ 2-x-5)
Answer by tutorcecilia(2144)

If the problem is (3x-5)+ (2-x-5):
3x-5+2-x-5 [Combine like-terms]
2x-8 [Set the expression equal to zero]
2x-8=0 [Solve for x]
2x-8+8=0+8
2x=8
2x/2=8/2
x=4
.
.
Check by plugging (x=4) back into the original equation:
(3x-5)+ (2-x-5)= 0 [Set the original equation equal to zero]
[(3)(4)-5)]+ (2-(4)-5)=0 [Simplify]
12-5+2-4-5=0
0=0 [Checks out]